When the sum of 467 and four times a positive number is subtracted from the square of the number, the result is 154. Find the number.

Answer by ankor@dixie-net.com(22740)

When the sum of 467 and four times a positive number is subtracted from the square of the number, the result is 154.
:
Write an equation for what is says
n^2 - (4n+467) = 154
n^2 - 4n - 467 = 154
n^2 - 4n - 467 - 154 = 0
A quadratic equation
n^2 - 4n - 621 = 0
The solution should be an integer, this will factor,
(n-27)(n+23) = 0
positive solution
n = 27
:
Check and see if that flies in the original equation.